OpenStudy (cutiecomittee123):

x^2+4y=100 4y-x^2=-20 How in the world so I solve this system of equations please help fast! Do I use substitution? I don't clearly see how I would solve for x or y in either equation.

OpenStudy (anonymous):

you can use substitution because x^2 on the first equation is positive and negative on the second equation. They will cancel out

OpenStudy (anonymous):

*as i was, the 4y's will cancel out

OpenStudy (cutiecomittee123):

so add them together

OpenStudy (cutiecomittee123):

I get y=10

OpenStudy (anonymous):

yup x^2-(-x^2)=2x^2 and then 100-(-20)=120 2x^2=120 120/2=60 x^2=60 and then take the square root

OpenStudy (cutiecomittee123):

yeah so square root of 60 = x

OpenStudy (anonymous):

exactly, and whatever x equals you plug it into to one of the equations to find y

OpenStudy (cutiecomittee123):

I plugged it into the second equation and it worked. So does that mean there is only one solution?

OpenStudy (anonymous):

no the number you plugged in is the answer for x. The answer you got after you plugged it in is y

OpenStudy (anonymous):

remember, there is an x and a y so there has to be at least 2 answers, one for each letter

OpenStudy (cutiecomittee123):

Yeah so (sqrt60, 10) is the only solution?

OpenStudy (cutiecomittee123):

I got the answer wrong it was actually (6,4) :(((